OpenStudy (anonymous):

1/4 - 3/2 + 9 - 54 +.... find the sum of the series if it converges

OpenStudy (campbell_st):

the common ratio for the series is r = -6 a series converges when |r|<1

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (campbell_st):

the common ratio is found by \[\frac{T _{2}}{T _{1}} = \frac{T _{3}}{T _{2}} = \frac{T _{4}}{T _{3}} ......\]

OpenStudy (campbell_st):

if the ratio is the same for the first few terms... then it will be assumed to be the same for all

OpenStudy (anonymous):

what is T1, T2, T3, and T4 then?

OpenStudy (anonymous):

I mean value wise

OpenStudy (anonymous):

T_1 = 1/4 T_2=- 3/2 ...

OpenStudy (campbell_st):

\[T _{1} = \frac{1}{4}, T _{2} = \frac{-3}{2}\]

OpenStudy (anonymous):

The sum of the series or a GP is \[S_n = \frac{a(r^n - 1)}{r-1}\] a-->1st term r--->common ration Substitute the values and you will get \[S_n = \frac{\frac{1}{4}({(-6)}^n - 1)}{-6-1}\] \[S_n = \frac{\frac{1}{4}({(-6)}^n - 1)}{-7}\] \[S_n = \frac{{((-6)}^n - 1)}{-28}\]

OpenStudy (anonymous):

*of

OpenStudy (anonymous):

Now you can find out whether it converges or not

OpenStudy (anonymous):

the common ratio for the series is r = -6 a series converges when |r|<1

OpenStudy (anonymous):

As Campbell said. So no need of doing all this. Just determine the r and find out. The above I did was just for explaining

OpenStudy (anonymous):

thanks